# Simultaneous Calibration of Noise Covariance and Kinematics for State Estimation of Legged Robots via Bi-level Optimization

### v2

@@ I Introduction
State estimation is fundamental to the autonomy of robotic systems, providing the basis for planning and control. Robots must infer their states–such as position, velocity, and orientation-from noisy sensor data, often under challenging real-world conditions. Classical approaches such as the Kalman filter and its extensions [kalman1960, bai2023state] remain central, and have been successfully deployed in applications ranging from ground vehicles [thrun2005probabilistic] and aerial robots [mohamed2021momentum] to quadrupeds and humanoids [bloesch2013state, rotella2014state, he2025invariant, HLIP]. More recently, factor graph–based methods and smoothing techniques [kaess2012isam2, dellaert2017factor] have enabled large-scale estimation for simultaneous localization and mapping (SLAM) [cadena2017past] and visual–inertial odometry [mur2017orb, usenko2019visual].
A persistent challenge in these methods is the specification of noise covariance matrices. The process noise covariance captures uncertainty in the dynamics and actuation, while the measurement noise covariance characterizes sensor error models. In practice, both are difficult to obtain: manufacturer datasheets provide only partial information, while individual sensor calibration and physical system identification are both expensive and task-dependent. As a result, practitioners often rely on manual tuning [he2025invariant, hartley2020contact, kang24] - a heuristic, time-consuming process that can produce estimators that are either sub-optimal or inconsistent [thrun2005probabilistic] over out-of-distribution tasks. This problem is particularly acute for highly dynamic robots such as quadrupeds and humanoids, whose complex, high-dimensional hybrid dynamics and kinematics, and multi-modal sensor nonlinearities amplify the challenges to obtain precise state estimates [rotella2014state, hartley2020contact, he2025invariant, HLIP, kang24], especially under kinematic uncertainties.
@@
Several lines of work have been proposed to address this issue. Adaptive filtering methods [9044358] adjust covariances online based on innovation statistics, while expectation–maximization (EM) frameworks [em_mhe_2017, 10472604, yap2008simultaneous] iteratively refine noise parameters to maximize likelihood [6630924, 6482657]. Learning sensor parameters or noises parameters [wong2020variational, qadri2024learning, yap2008simultaneous] is another common approach. However, these methods often either assume no availability of ground truth data or only focus on decoupled calibrations. More importantly, their treatments have not been examined on modern legged robots that have multimodal sensor suites and complex dynamics.
In this work, we propose a bi-level optimization framework for simultaneous covariance and kinematics calibration for robotic state estimation. At the upper level, we treat covariance matrices and uncertain kinematics as optimization variables. At the lower level, we solve a full-information estimator using the Maximum a Posteriori (MAP) formulation given the current covariance estimates and kinematics information. By differentiating through the estimator [amos2017optnet, icml2022implicit], we directly optimize parameters to be calibrated with respect to trajectory-level objectives of minimizing the errors between the estimates and the measurable ground truth. This approach eliminates manual tuning, enforces physical consistency through structured parameterization and constraints, and generalizes across sensor modalities and robots.
@@ II Related Work
Legged Robotic State Estimation: State estimation for legged robots has been extensively studied due to the challenges posed by intermittent ground contacts, actuator uncertainties, and highly dynamic motions. Early approaches utilize Extended Kalman Filters (EKFs) to fuse proprioceptive sensing - Inertial Measurement Unit (IMU), contact sensors, and joint encoders - with kinematic constraints to achieve drift-reduced estimates on quadrupedal [bloesch2013state], bipedal [teng2021legged] and humanoid robots [rotella2014state]. Recent focuses of using EKFs on legged robots employ the invariant properties of the SE(3) manifold, and thus yield more accurate estimates [he2025invariant, hartley2020contact]. Another line of work leveraged factor graphs [dellaert2017factor, yang2024real] and smoothing or windowed formulations such as Moving Horizon Estimation (MHE) [kang24] to integrate proprioceptive and exteroceptive sensing data at scale. Despite these advances, most approaches rely on fixed or manually tuned noise covariances, limiting the accuracy of estimates, especially in the presence of kinematic uncertainty. Our work addresses this gap by introducing a bi-level optimization framework that calibrates covariance matrices and uncertain kinematics simultaneously in a principled manner.
@@
Sensor Calibration: Sensor calibration and noise characterization have also been studied extensively in robotics, since the accuracy of estimation strongly depends on the quality of sensor models. Traditional approaches focus on extrinsic and intrinsic calibration of individual sensors, such as cameras [rehder2016extending], LiDARs [li2023automatic], and IMUs [fu2021high]. At the theoretical level, several works address the identification of noise statistics. For instance, expectation–maximization (EM) algorithms iteratively refine noise models to maximize likelihood [em_mhe_2017, 10472604, yap2008simultaneous]. Recent efforts explored data-driven estimation of uncertainty, including machine learning to optimize covariances from raw sensory inputs [qadri2024learning, wang2023neural]. Despite these advances, most existing techniques commonly are tailored to specific sensors rather than complex legged robots; they often assume restrictive statistical models, or require careful initialization. In contrast, our bi-level optimization jointly calibrates both process and measurement noise with uncertain kinematics across diverse and complex legged robotic platforms.
@@
System Identification: System identification (SysID) has long been a cornerstone of robotics, enabling accurate modeling of dynamics, kinematics, and actuation. Classical techniques for parameter estimation of rigid-body kinematics or dynamics exploit regressions of the geometric or Lagrangian models [leboutet2021inertial, 7987066, khorshidi2025physically] subject to physical constraints. Gaussian processes or neural networks approximate residuals not captured by physics-based models [hwangbo2019learning] are commonly used in calibration. Recent works [yang2022online, burgul2024online] focus on online kinematics calibration via state estimators. Despite significant progress, SysID typically treats noise statistics, model parameters, and estimator design separately. Our bi-level optimization approach bridges this gap by jointly calibrating positive definite covariance matrices and kinematic parameters, highlighting the natural synergy between system identification and state estimation.
@@ III Preliminaries
We first introduce the models of legged state estimation that are widely used in the literature [rotella2014state, bloesch2013state, hartley2020contact, kang24, he2025invariant]. Then, we will elaborate the preliminaries of the mathematics on formulating state estimation as an optimization problem that we will employ in our calibration work. Throughout, $\mathrm{W}$ denotes the world frame and $\mathrm{B}$ the body (base) frame.
@@ III-A1 Process Model
We consider a discrete-time nonlinear process model for state $\mathbf{x}_{k}$ and input $\mathbf{u}_{k}$ at time step $k=0,\ldots,T-1$ ($T$ denotes the length of the discrete horizon):
$\mathbf{x}_{k+1}=\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{w}_{k},\quad\mathbf{w}_{k}\sim\mathcal{N}(0,Q_{k}).$ (1)
The robot-centric state stacks the pose and the velocity of the floating-base, foot positions, and IMU biases: $\mathbf{x}_{k}=\begin{bmatrix}p^{\top}&v^{\top}&q^{\top}&p_{\text{foot},k}^{\top}&b_{a}^{\top}&b_{\omega}^{\top}\end{bmatrix}^{\top}$, where $p$ and $v$ are the linear position and velocity of the torso in $\mathrm{W}$, $q$ denotes the unit quaternion that parameterizes the rotation matrix $\mathbf{R}_{\mathrm{WB}}\in SO(3)$ of the torso w.r.t. $\mathrm{W}$, $p_{\text{foot},k}$ contains all the foot positions that are described in $\mathrm{W}$, and $b_{a},b_{\omega}\in\mathbb{R}^{3}$ are the biases terms of the IMU in $\mathrm{B}$. For later use, we write the state as a product-manifold element $\mathbf{x}_{k}=(\mathbf{T}_{k},\mathbf{e}_{k},b_{a,k},b_{\omega,k})\in\mathrm{SE}(3)\times\mathbb{R}^{n_{\mathrm{eucl}}}$, where $\mathbf{T}_{k}\in\mathrm{SE}(3)$ encodes the floating-base transformation with rotation parameterized by quaternion $q_{k}$, and $\mathbf{e}_{k},$ stacks the Euclidean sub-states, i.e., the base velocity $v_{k}$ and foot positions, $n_{\mathrm{eucl}}$ denotes the dimension of the Euclidean substate $\mathbf{e}_{k}$. The IMU measures the body acceleration $a$ and angular velocity $\omega$ with the standard additive biases and noises:
$\tilde{a}$ $=a+b_{a}+\delta_{a},$ $\delta_{a}$ $\sim\mathcal{N}(0,Q_{a}),$ (2)
$\tilde{\omega}$ $=\omega+b_{\omega}+\delta_{\omega},\quad$ $\delta_{\omega}$ $\sim\mathcal{N}(0,Q_{\omega}),$
and the biases have random walk dynamics:
$b_{a}^{+}$ $=b_{a}+\delta_{b_{a}},$ $\delta_{b_{a}}$ $\sim\mathcal{N}(0,Q_{b_{a}}),$ (3)
$b_{\omega}^{+}$ $=b_{\omega}+\delta_{b_{\omega}},$ $\delta_{b_{\omega}}$ $\sim\mathcal{N}(0,Q_{b_{\omega}}).$
Using these IMU measurements for state propagation, the discrete-time updates in (1) are:
$p^{+}$ $=p+v\,\Delta t+\tfrac{1}{2}\!\left(\mathbf{R}_{\mathrm{WB}}(\tilde{a}-b_{a})+g\right)\!\Delta t^{2}\;+\;\delta_{p},$
$v^{+}$ $=v+\left(\mathbf{R}_{\mathrm{WB}}(\tilde{a}-b_{a})+\mathbf{g}\right)\!\Delta t\;+\;\delta_{v},$
@@
$q^{+}$ $=\zeta\!\left((\tilde{\omega}-b_{\omega})\,\Delta t\right)\otimes q\;+\;\delta_{q},$
$p_{\text{foot}}^{+}$ $=p_{\text{foot}}+\delta_{\text{foot}},\quad b_{a}^{+}=b_{a}+\delta_{b_{a}},\quad b_{\omega}^{+}=b_{\omega}+\delta_{b_{\omega}},$
where $\otimes$ denotes quaternion multiplication and $\zeta(\cdot)$ maps a small rotation vector to the corresponding unit quaternion, $g$ is gravitational acceleration in $\mathrm{W}$. The process noises $\delta_{(\cdot)}$ are modeled as zero-mean Gaussian $\mathcal{N}(0,Q_{(\cdot)})$, which we collect into a stacked random vector: $\mathbf{w}\!=\![\delta_{p}^{\top},\delta^{\top}_{v},\delta^{\top}_{q},\delta^{\top}_{\text{foot}},\delta^{\top}_{b_{a}},\delta^{\top}_{b_{\omega}}]^{\top}\!\sim\mathcal{N}\!\big(0,Q_{\mathbf{w}}\big)$, and $Q_{\mathbf{w}}=\operatorname{blkdiag}\!\big(Q_{p},Q_{a},Q_{\omega},Q_{\mathrm{foot}},Q_{b_{a}},Q_{b_{\omega}}\big)$. In practice, the random-walk covariances are scaled proportionally to the sampling period $\Delta t$.
@@ III-A2 Measurement Model
The multi-sensor measurements $\mathbf{y}_{k}$ are functions of the state and input with Gaussian noise:
$\mathbf{y}_{k}=\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{n}_{k},\qquad\mathbf{n}_{k}\sim\mathcal{N}(0,R_{k}).$ (4)
The observation model $\mathbf{h}(\cdot)$ aggregates all available sensors. Joint encoders provide the measurements of the joint angles $\tilde{\alpha}$ and angular velocities $\dot{\tilde{\alpha}}$, which are assumed with Gaussian noises: $\tilde{\alpha}=\alpha+\delta_{\alpha}$ and $\dot{\tilde{\alpha}}=\dot{\alpha}+\delta_{\dot{\alpha}}$, where $\delta_{\alpha}\!\sim\!\mathcal{N}(0,R_{\alpha})$ and $\delta_{\dot{\alpha}}\!\sim\!\mathcal{N}(0,R_{\dot{\alpha}})$. Given leg kinematics, forward kinematics $fk(\cdot)$ and Jacobian $J(\cdot)$ yield the measurements of the foot positions and velocities w.r.t. the torso:
[…]
$y_{v}$ $=v+\delta_{y_{v}},$
[…]
where $\cdot$$_{\times}$ denotes skew-symmetric map, $\delta_{\mathrm{pf}}\!\sim\!\mathcal{N}(0,R_{\mathrm{pf}})$ and $\delta_{\mathrm{vf}}\!\sim\!\mathcal{N}(0,R_{\mathrm{vf}})$ encapsulates kinematics errors and encoder/gyroscope noises [teng2021legged]. Here $\tilde{(\cdot)}$ denotes a raw sensor measurement.
For simplicity, here we consider a single-leg situation; the multi-leg case follows by block-diagonal stacking. Let the proprioceptive noise $\delta z\!=\![\delta\alpha^{\top}\ \delta\dot{\alpha}^{\top}\ \delta\omega^{\top}]^{\top}$, and thus,
$\delta z\sim\mathcal{N}\!\left(0,\;R_{z}\right),R_{z}=\mathrm{blkdiag}\big(R_{\alpha},R_{\dot{\alpha}},Q_{\omega}\big)\in\mathbb{S}_{++}^{9},$ (6)
where the operator $\mathrm{blkdiag}(\cdot)$ forms a block-diagonal matrix from its arguments and $\mathbb{S}_{++}^{n}$ denotes the cone of $n\times n$ symmetric positive-definite matrices.. A first-order perturbation of the foot channels in (5) gives $\delta y_{\mathrm{pf}}\;=\;G_{p}\,\delta z$, $\delta y_{\mathrm{vf}}\;=\;G_{v}\delta z$, with Jacobian matrices as:

### v3

@@ I Introduction
State estimation is fundamental to the autonomy of robotic systems, providing the basis for planning and control. Robots must infer their states–such as position, velocity, and orientation-from noisy sensor data, often under challenging real-world conditions. Classical approaches such as the Kalman filter and its extensions [21, 3] remain central, and have been successfully deployed in applications ranging from ground vehicles [32] and aerial robots [26] to quadrupeds and humanoids [5, 30, 16, 38]. More recently, factor graph–based methods and smoothing techniques [20, 10] have enabled large-scale estimation for simultaneous localization and mapping (SLAM) [7] and visual–inertial odometry [27, 33].
A persistent challenge in these methods is the specification of noise covariance matrices. The process noise covariance captures uncertainty in the dynamics and actuation, while the measurement noise covariance characterizes sensor error models. In practice, both are difficult to obtain: manufacturer datasheets provide only partial information, while individual sensor calibration and physical system identification are both expensive and task-dependent. As a result, practitioners often rely on manual tuning [16, 15, 22] - a heuristic, time-consuming process that can produce estimators that are either sub-optimal or inconsistent [32] over out-of-distribution tasks. This problem is particularly acute for highly dynamic robots such as quadrupeds and humanoids, whose complex, high-dimensional hybrid dynamics and kinematics, and multi-modal sensor nonlinearities amplify the challenges to obtain precise state estimates [30, 15, 16, 38, 22], especially under kinematic uncertainties.
@@
Several lines of work have been proposed to address this issue. Adaptive filtering methods [42] adjust covariances online based on innovation statistics, while expectation–maximization (EM) frameworks [12, 14, 41] iteratively refine noise parameters to maximize likelihood [4, 9]. Learning sensor parameters or noises parameters [37, 28, 41] is another common approach. However, these methods often either assume no availability of ground truth data or only focus on decoupled calibrations. More importantly, their treatments have not been examined on modern legged robots that have multimodal sensor suites and complex dynamics.
In this work, we propose a bi-level optimization framework for simultaneous covariance and kinematics calibration for robotic state estimation. At the upper level, we treat covariance matrices and uncertain kinematics as optimization variables. At the lower level, we solve a full-information estimator using the Maximum a Posteriori (MAP) formulation given the current covariance estimates and kinematics information. By differentiating through the estimator [1, 13], we directly optimize parameters to be calibrated with respect to trajectory-level objectives of minimizing the errors between the estimates and the measurable ground truth. This approach eliminates manual tuning, enforces physical consistency through structured parameterization and constraints, and generalizes across sensor modalities and robots.
@@ II Related Work
Legged Robotic State Estimation: State estimation for legged robots has been extensively studied due to the challenges posed by intermittent ground contacts, actuator uncertainties, and highly dynamic motions. Early approaches utilize Extended Kalman Filters (EKFs) to fuse proprioceptive sensing - Inertial Measurement Unit (IMU), contact sensors, and joint encoders - with kinematic constraints to achieve drift-reduced estimates on quadrupedal [5], bipedal [31] and humanoid robots [30]. Recent focuses of using EKFs on legged robots employ the invariant properties of the SE(3) manifold, and thus yield more accurate estimates [16, 15]. Another line of work leveraged factor graphs [10, 40] and smoothing or windowed formulations such as Moving Horizon Estimation (MHE) [22] to integrate proprioceptive and exteroceptive sensing data at scale. Despite these advances, most approaches rely on fixed or manually tuned noise covariances, limiting the accuracy of estimates, especially in the presence of kinematic uncertainty. Our work addresses this gap by introducing a bi-level optimization framework that calibrates covariance matrices and uncertain kinematics simultaneously in a principled manner.
@@
Sensor Calibration: Sensor calibration and noise characterization have also been studied extensively in robotics, since the accuracy of estimation strongly depends on the quality of sensor models. Traditional approaches focus on extrinsic and intrinsic calibration of individual sensors, such as cameras [29], LiDARs [25], and IMUs [11]. At the theoretical level, several works address the identification of noise statistics. For instance, expectation–maximization (EM) algorithms iteratively refine noise models to maximize likelihood [12, 14, 41]. Recent efforts explored data-driven estimation of uncertainty, including machine learning to optimize covariances from raw sensory inputs [28, 35]. Despite these advances, most existing techniques commonly are tailored to specific sensors rather than complex legged robots; they often assume restrictive statistical models, or require careful initialization. In contrast, our bi-level optimization jointly calibrates both process and measurement noise with uncertain kinematics across diverse and complex legged robotic platforms.
@@
System Identification: System identification (SysID) has long been a cornerstone of robotics, enabling accurate modeling of dynamics, kinematics, and actuation. Classical techniques for parameter estimation of rigid-body kinematics or dynamics exploit regressions of the geometric or Lagrangian models [24, 36, 23] subject to physical constraints. Gaussian processes or neural networks approximate residuals not captured by physics-based models [18] are commonly used in calibration. Recent works [39, 6] focus on online kinematics calibration via state estimators. Despite significant progress, SysID typically treats noise statistics, model parameters, and estimator design separately. Our bi-level optimization approach bridges this gap by jointly calibrating positive definite covariance matrices and kinematic parameters, highlighting the natural synergy between system identification and state estimation.
@@ III Preliminaries
We first introduce the models of legged state estimation that are widely used in the literature [30, 5, 15, 22, 16]. Then, we will elaborate the preliminaries of the mathematics on formulating state estimation as an optimization problem that we will employ in our calibration work.
@@ III-A1 Process Model
We define the state as the composition of the floating-base kinematics, the absolute positions of the footholds, and the IMU biases in a discrete-time model:
$\mathbf{x}_{k}:=\left(\mathbf{p}_{k},\mathbf{v}_{k},\mathbf{q}_{k},\mathbf{p}_{\text{foot},k},\mathbf{b}_{a,k},\mathbf{b}_{\omega,k}\right).\vskip-1.0pt$ (1)
Here $\mathbf{p}_{k}$ and $\mathbf{v}_{k}$ denote the position and velocity of the main body expressed in inertial frame $\mathrm{W}$. The orientation is represented by a unit quaternion $\mathbf{q}_{k}$ that parameterizes the rotation from $\mathrm{W}$ to the body frame $\mathrm{B}$. To incorporate leg kinematics, the state is augmented with the absolute positions of the footholds $\mathbf{p}_{\text{foot},k}$ in $\mathrm{W}$, the accelerometer and gyroscope biases $\mathbf{b}_{a,k},\mathbf{b}_{\omega,k}$ expressed in $\mathrm{B}$. Formally, the state $\mathbf{x}_{k}$ resides on the product manifold $\mathcal{G}:=\mathrm{SE}(3)\times\mathbb{R}^{n_{\text{aug}}}$, where the pose $(\mathbf{p}_{k},\mathbf{q}_{k})$ constitutes the $\mathrm{SE}(3)$ component and the remaining Euclidean sub-states are stacked in $\mathbb{R}^{n_{\text{aug}}}$.
The IMU provides measurements of the proper acceleration $\tilde{\mathbf{a}}_{k}$ and the angular rate $\tilde{\bm{\omega}}_{k}$ of the robot’s main body[22]. These quantities are modeled as the true values corrupted by slowly varying biases and additive Gaussian white noise:
$\tilde{\mathbf{a}}_{k}=\mathbf{a}_{k}+\mathbf{b}_{a,k}+\mathbf{w}_{a,k},\qquad\tilde{\bm{\omega}}_{k}=\bm{\omega}_{k}+\mathbf{b}_{\omega,k}+\mathbf{w}_{\omega,k},\vskip-3.0pt$
where $\mathbf{a}_{k}$ is the absolute acceleration rotated into frame $\mathrm{B}$ minus the gravitational vector, and $\mathbf{w}_{(\cdot)}$ represents the discrete Gaussian noise terms. Assuming a zero-order hold (ZOH) for the bias-corrected IMU measurements[5]. The discrete-time prediction equations are formulated as follows:
$\mathbf{p}_{k+1}$ $=\mathbf{p}_{k}\!+\!\mathbf{v}_{k}\Delta t\!+\!\tfrac{1}{2}\left(\mathbf{C}(\mathbf{q}_{k})(\tilde{\mathbf{a}}_{k}\!-\!\mathbf{b}_{a,k})\!+\!\mathbf{g}\right)\Delta t^{2}\!+\!\mathbf{w}_{p},$
$\mathbf{v}_{k+1}$ $=\mathbf{v}_{k}+\left(\mathbf{C}(\mathbf{q}_{k})(\tilde{\mathbf{a}}_{k}-\mathbf{b}_{a,k})+\mathbf{g}\right)\Delta t+\mathbf{w}_{v},$
@@
$\mathbf{q}_{k+1}$ $=\zeta\left((\tilde{\bm{\omega}}_{k}-\mathbf{b}_{\omega,k})\Delta t\right)\otimes\mathbf{q}_{k}+\mathbf{w}_{q},$
$\mathbf{p}_{\text{foot},k+1}$ $=\mathbf{p}_{\text{foot},k}+\mathbf{w}_{\text{foot}},\qquad\mathbf{b}_{(\cdot),k+1}=\mathbf{b}_{(\cdot),k}+\mathbf{w}_{b_{(\cdot)}}.\vskip-3.0pt$
where $\mathbf{C}(\mathbf{q}_{(\cdot)})$ extracts the rotation matrix from quaternion, and $\mathbf{g}$ is the gravity vector in $\mathrm{W}$, $\otimes$ denotes the quaternion multiplication operator and $\zeta(\cdot)$ represents the map from a rotation vector to its corresponding unit quaternion[5, 22]. Although footholds are assumed stationary, a random walk is introduced to accommodate potential slippage, with IMU biases likewise modeled as Brownian motions. The aggregated process noise $\mathbf{w}_{k}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{Q}_{k})$ follows the uncertainty modeling architecture detailed in Section IV.B of [22].
@@ III-A2 Measurement Model
We formulate the measurement model for a generic leg by relating the derived kinematics $\tilde{\mathbf{y}}_{k}$ to the state-predicted observations $\mathbf{h}(\mathbf{x}_{k})$. The full multi-leg model follows by stacking individual residuals, implying a block-diagonal covariance structure under the assumption of independent sensor noise across legs. This relationship accounts for the mapping of raw sensor uncertainty into the task space via an additive effective noise term $\mathbf{n}_{k}$ [5]:
$\tilde{\mathbf{y}}_{k}=\mathbf{h}(\mathbf{x}_{k})+\mathbf{n}_{k},\quad\mathbf{n}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{k}).\vskip-4.0pt$
Kinematic Measurements: Joint encoders and the IMU gyroscope provide raw measurements $\tilde{\mathbf{z}}\!:=\![\tilde{\bm{\alpha}}\!^{\top}\!,\!\dot{\tilde{\bm{\alpha}}}\!^{\top}\!,\!\tilde{\bm{\omega}}\!^{\top}$$^{\top}$ consisting of joint angles/velocities $\tilde{\bm{\alpha}}/\dot{\tilde{\bm{\alpha}}}$ and body angular velocities $\tilde{\bm{\omega}}$. We model these raw inputs as the true values corrupted by additive Gaussian noise $\bm{\eta}\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma}_{\eta})$, where $\mathbf{\Sigma}_{\eta}=\mathrm{blkdiag}(\mathbf{R}_{\alpha},\mathbf{R}_{\dot{\alpha}},\mathbf{Q}_{\omega})$. Leveraging the forward kinematics $\mathrm{fk}(\cdot)$ and the analytical leg Jacobian $\mathbf{J}(\cdot)$, $\mathbf{h}(\mathbf{x}_{k})$ and $\tilde{\mathbf{y}}_{k}$ derived from raw sensor inputs are defined as follows:
$\mathbf{h}_{p}(\mathbf{x}_{k})$ $=\mathbf{C}(\mathbf{q}_{k})^{\top}\left(\mathbf{p}_{\text{foot},k}-\mathbf{p}_{k}\right),$ (2)
$\tilde{\mathbf{y}}_{p,k}$ $=\mathrm{fk}(\tilde{\bm{\alpha}}_{k}),$
$\mathbf{h}_{v}(\mathbf{x}_{k})$ $=-\mathbf{C}(\mathbf{q}_{k})^{\top}\mathbf{v}_{k},$
$\tilde{\mathbf{y}}_{v,k}$ $=\mathbf{J}(\tilde{\bm{\alpha}}_{k})\dot{\tilde{\bm{\alpha}}}_{k}+[\tilde{\bm{\omega}}_{k}-\mathbf{b}_{\omega,k}]_{\times}\mathrm{fk}(\tilde{\bm{\alpha}}_{k}).$
where $\cdot$$_{\times}$ denotes the skew-symmetric operator, $\mathbf{h}_{v}$ models negative body velocity under the static contact assumption, while $\tilde{\mathbf{y}}_{v}$ captures the actual proprioceptive measurement.
Uncertainty Modeling: A critical aspect of our formulation is the analytic propagation of sensor noise into the task space. Following the linearization approach in, the covariance matrices $\mathbf{R}_{p,k}$ and $\mathbf{R}_{v,k}$ for the position and velocity residuals are derived via a first-order expansion of the kinematic functions w.r.t. the noise vector $\bm{\eta}$: $\mathbf{R}_{k}=\mathbf{G}_{k}\mathbf{\Sigma}_{\eta}\mathbf{G}_{k}^{\top}+\mathbf{R}_{\text{add}},$ where $\mathbf{R}_{\text{add}}$ accounts for unmodeled kinematic calibration errors, and the mapping matrix $\mathbf{G}_{k}\!:=\partial\tilde{\mathbf{y}}_{k}/\partial\tilde{\mathbf{z}}_{k}$ is the Jacobian of the kinematic measurements $\tilde{\mathbf{y}}_{k}\!:=\![\tilde{\mathbf{y}}_{p,k}\!^{\top}\!,\!\tilde{\mathbf{y}}_{v,k}\!^{\top}]^{\top}$ w.r.t $\tilde{\mathbf{z}}_{k}$: